3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 3(x-4) = 21

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3(x-4) = 21

3x – 12 = 21

3x = 21 + 12

3x = 33

x = 11

Thanks